# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v4

@@ E.2.2 Analysis
For estimator $\hat{\mu}_{i}$ that Algorithm 8 builds, it holds that:
$\Pr\left[\left|\hat{\mu}_{i}-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]\right|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\geq 1-\frac{1}{T^{2}},$
@@
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right),$
$R_{s,i}$ is the $s$-th reward sample from $\textsf{Bern}(r_{i}q_{t})$ when arm $i$ is pulled in line $2$ and $q_{0,i}$ is the value of starting state before starting pull arm $i$.
@@
Let $q_{0,i}$ be the value of the state before we start pulling arm $i$ for $(T/K)^{2/3}$ rounds. Then, after $(T/K)^{2/3}$ pulls of this arm the state becomes:
$q_{(T/K)^{2/3}+1,i}$ $=(1-\lambda)^{(T/K)^{2/3}}q_{0,i}+\lambda\sum_{s=0}^{T^{2/3}-1}(1-\lambda)^{s}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}q_{0,i}+\lambda\cdot\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{1-(1-\lambda)}b_{i}$
@@
Then, for the next $(T/K)^{2/3}$ rounds the algorithm pulls arm $i$ and samples reward from
$R_{s,i}\sim\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+s,i})$ for $s\in[(T/K)^{2/3}]$. For each $s\in[(T/K)^{2/3}]$, the state is:
$q_{(T/K)^{2/3}+s,i}$ $=(1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}.$
The expected reward collected for rounds at each $s\in[(T/K)^{2/3}]$ is
$\mathbb{E}[R_{s,i}]=\mathbb{E}\left[R_{s,i}\sim\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+s,i})\right]$ $=r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
Summing for all $s\in[K]$:
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
Using union bound over all arms we have Using Hoeffding’s inequality on $\hat{\mu}_{i}$ we get the following result:
$\Pr\left[\left|\hat{\mu}_{i}-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]\right|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\geq 1-\frac{1}{T^{2}},$ (33)
Using union bound over all arms we have the result. ∎
@@
Fix a pair $(i,j)$, where $i,j\in[K]$. Then, for the estimators $\hat{\mu}_{i,j}$ that Algorithm 8 builds it holds that:
$\Pr\left[\left|\hat{\mu}_{i,j}-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{T^{2/3}-1}R_{s,i}^{\text{ALT}}\right]\right|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$
@@
$\mathbb{E}\left[\sum_{s=0}^{T^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+W_{i,j}(\lambda),$
$q_{0,i}^{\text{ALT}}$ is the value of state before we start alternating arms and $R_{s,i}^{\text{ALT}}$ is the reward sampled during the alternation between arms $i,j$.
@@
Algorithm 8 first alternates between arm $i$ and arm $j$ for $(T/K)^{2/3}$ rounds. After this many rounds, the state becomes (see Lemma E.3):
$q_{(T/K)^{2/3}+1,i}^{\text{ALT}}$ $=(1-\lambda)^{(T/K)^{2/3}}q_{0,i}^{\text{ALT}}+\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{2-\lambda}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+W_{i,j}(\lambda)$
Next, for another $2(T/K)^{2/3}$ rounds the algorithm keeps alternating between the two arms and samples reward from $R_{s,i}^{\text{ALT}}\sim\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+2s+1,i}^{\text{ALT}})$, where
$q_{(T/K)^{2/3}+2s+1,i}^{\text{ALT}}$ $=(1-\lambda)^{(T/K)^{2/3}+2s}q_{0,i}^{\text{ALT}}+\frac{1-(1-\lambda)^{(T/K)^{2/3}+2s}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{(T/K)^{2/3}+2s}}{2-\lambda}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+W_{i,j}(\lambda)$
Using this expression for the state, the expected reward obtained is equal:
$\mathbb{E}\left[\sum_{s=0}^{T^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+r_{i}W_{i,j}(\lambda)$
Finally, using Hoeffding’s inequality as in Lemma E.4 on $\hat{\mu}_{i,j}$, we get the result. ∎
Taking a union bound over all $i\in[K]$ and the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ they correspond to, we get that Lemmas E.4 and E.5 will hold for all $i,j\in[K]^{2}$ with probability at least $2/T$ as $K<<T$.
Next, in Lemma E.6 we bound the regret of Algorithm 8 for the case where $\lambda$ is unknown and small. More specifically, we are going to prove that Algorithm 8 after building the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ with the properties outlined above is going to call EXP3.P. The regret then follows from the results of Section 4.
@@
If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$, then Algorithm 8 incurs regret:
@@
At a high level, our proof proceeds as follows: (1) If the true $\lambda$ is near 00, then the state changes very little throughout the $T$ rounds; (2) the difference between $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ from Algorithm 8 is very small too; (3) Algorithm 8 calls EXP3.P.
Originally (i.e., in round $1$), the state is equal to $q_{1}=q_{0}=1$. Then, Algorithm 8 pulls all arms $i\in[K]$ for $K^{1/3}T^{2/3}$ rounds. If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$, the values of the states $q_{t}$ during the $K^{1/3}T^{2/3}$ rounds of building the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ will be close to each other. To see this, notice that: $1\geq(1-\lambda)^{K^{1/3}T^{2/3}}\geq(1-1/T)^{K^{1/3}T^{2/3}}>1-\frac{1}{T^{1/3}}$, which confirms that the state does not change significantly during $K^{1/3}T^{2/3}$ rounds. After a union bound over all arms $i\in[K]$ for both Lemmas E.4 and E.5, we get that with probability $1-2/T$:
$\hat{\mu}_{i}-\hat{\mu}_{i,j}\leq 2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$
and also:
$\hat{\mu}_{i}-\hat{\mu}_{i,j}\geq-2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$
Thus, to bound $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|$ we want to bound the difference $\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$:
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
$\quad\quad-\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+r_{i}W_{i,j}(\lambda)$
$=\sum_{s=0}^{(T/K)^{2/3}-1}\underbrace{r_{i}\left((1-\lambda)^{(T/K)^{2/3}+s}q_{0,i}-(1-\lambda)^{(T/K)^{2/3}+2s}q_{0,i}^{ALT}\right)}_{A_{1}}$
$\quad\quad+\underbrace{r_{i}b_{i}\left(1-(1-\lambda)^{(T/K)^{2/3}+s}\right)}_{A_{2}}+\underbrace{r_{i}W_{i,j}\left(1-(1-\lambda)^{(T/K)^{2/3}+2s}\right)}_{A_{3}}$
Now remember that $(1-\lambda)^{K^{1/3}T^{2/3}}\geq 1-\frac{1}{(T/K)^{1/3}}$, where $K^{1/3}T^{2/3}$ are the rounds spent for building estimators $\hat{\mu}_{i}$s and $\hat{\mu}_{i,j}$s. Thus, the value of the state $q_{t}$ during these rounds takes values: $1-\frac{1}{(T/K)^{1/3}}\leq q_{t}\leq 1$, which implies that $-\Theta(1/(T/K)^{1/3})\leq A_{1}\leq\Theta(1/(T/K)^{1/3})$, $0\leq A_{2}\leq(1/(T/K)^{1/3})$ and $0\leq A_{3}\leq(1/(T/K)^{1/3})$. Hence,
$\frac{1}{(T/K)^{2/3}}\left|\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]\right|$ $\leq\Theta(1/(T/K)^{1/3}),$
and $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq 5\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ with probability $1-2/T$. Conditioning on this event, Algorithm 8 will call EXP3.P (Algorithm 6). From Section 4, we get that the regret in this case is:
$R_{\textsc{DES}}(T)\leq\underbrace{2K^{1/3}T^{2/3}}_{\text{exploration samples}}+\underbrace{\mathcal{O}\left(\sqrt{KT\log T}\right)+\left(1-(1-\lambda)^{T}\right)\textsf{OPT}}_{\text{$R_{\textsc{DES}}$ of $\textsf{EXP3.P}$}}.$
Accounting for the regret picked up in the case where the event does not hold only adds an extra constant to the regret. This concludes our proof. ∎
Next, we bound the regret for $\lambda>\omega((K/T)^{2/3})$. To do this, we divide the analysis in two cases. In the first case (Lemma E.7), we have: $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(K^{1/3}T^{-1/3}\sqrt{\log T})$. Note that in this case, learning $\lambda$ (even approximately) is rather difficult. However, the difference between estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ is rather small because by assumption the arms are not that different from each other. In the second case (Lemma E.8), we have $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}>\omega(K^{1/3}T^{-1/3}\sqrt{\log T})$. For this case, Algorithm 8 is deployed in order to learn $\lambda$.
@@
If $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(\sqrt{\log T}/(T/K)^{1/3})$ and $\lambda>(K/T)^{2/3}$, then Algorithm 8 incurs regret $R_{\textsc{DES}}(T)=\widetilde{\mathcal{O}}\left(K^{1/3}T^{2/3}\right)$.

### v5

@@ 17             Play arm $j$
For each $N(\lambda)$ the estimate $\hat{\lambda}$ is between:
$1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}-\Omega(\delta)\leq\hat{\lambda}\leq 1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}+O(\delta),$
with probability at least $4\exp(-2M\delta^{2})$ under the assumption of $(A):|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|\geq o(\delta)$ and $|\mathbb{E}[\hat{r}_{i,i}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta\geq o(\delta)$. Where $\hat{r}_{i,i}$ is the estimator that algorithm builds of random variable $R_{i,i}\sim\textsf{Bern}\left(r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{i})+b_{i}\right)\right)$, $q_{0,i}$ is the value of state before we start sampling. Respectively, $\hat{r}_{i,b_{j}}$ is the estimator that algorithm builds of random variable $R_{i,b_{j}}\sim\textsf{Bern}\left(r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{j})+b_{j}\right)\right)$, $q_{0,b_{j}}$ is the value of state before we start sampling. And $\hat{r}_{i,j}$ is the estimator that algorithm builds of random variable $R_{i,j}\sim\textsf{Bern}\left(r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,j}-\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda})+\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right)\right)$ where $q_{0,j}$ is the value of state before we start sampling.
Proof.
For the estimator $\hat{r}_{i,b_{j}}$ we begin from a random state $q_{0,b_{j}}$ and then the we play arm $j$ for $\tilde{N}(\lambda)$ rounds so the state becomes from Lemma 3.5:
$q_{t}=(1-\lambda)^{\tilde{N}(\lambda)}(q_{0,b_{j}}-b_{j})+b_{j}.$
Using Hoeffding’s Inequality as in Lemma 3.8 we get:
$\Pr\left[\left|\hat{r}_{i,b_{j}}-r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,b_{j}}-b_{j})+b_{j}\right)\right|\geq\delta\right]\leq 2\exp\left(-2M\delta^{2}\right)$
Then, for $\hat{r}_{i,i}$ we fist play arm $j$ for $\tilde{N}(\lambda)$ rounds and the state equals to another random value $q_{0,i}$. Then, we play arm $i$ for $\tilde{N}(\lambda)$ rounds and the state becomes $q_{t}=(1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{i})+b_{i}$ and then we take a sample from $\textsf{Bern}(r_{i}q_{t})$. Thus, the expected value of $\mathbb{E}[\hat{r}_{i,i}]=r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{r}-b_{i})+b_{i}\right)$, and using Hoeffding’s inequality we get:
$\Pr\left[\left|\hat{r}_{i,j}-r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}(q_{0,i}-b_{i})+b_{j}\right)\right|\geq\delta\right]\leq 2\exp\left(-2M\delta^{2}\right)$
Then the same applies for $\hat{r}_{i,j}$:
$\Pr\left[\left|\hat{r}_{i,j}-r_{i}\cdot\left((1-\lambda)^{\tilde{N}(\lambda)}\left(q_{0,j}-\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right)+\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right)\right|\geq\delta\right]\leq 2\exp\left(-2M\delta^{2}\right)$
Thus, $\hat{\lambda}$ is upper bounded by:
$\hat{\lambda}$ $=\frac{\hat{r}_{i,b_{i}}+\hat{r}_{i,b_{j}}-2\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}$
$=1-\left|\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\right|\leq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|+2\delta}$ ( $\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\leq 0$ as ($\lambda\leq 1$))
$\leq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|}+O(\delta)$ (From assumption (A))
$=1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}+O(\delta)$
and lower bounded by
$\hat{\lambda}$ $=\frac{\hat{r}_{i,b_{i}}+\hat{r}_{i,b_{j}}-2\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}$
$=1-\left|\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\right|\geq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|+2\delta}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta}$ ( $\frac{\hat{r}_{i,b_{j}}-\hat{r}_{i,j}}{\hat{r}_{i,b_{i}}-\hat{r}_{i,j}}\leq 0$ as ($\lambda\leq 1$))
$=1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|}-\Omega(\delta)$ (From assumption $(A)$)
$=1+\frac{\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]}{\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]}-\Omega(\delta)$
These are satisfied with probability at least $4\exp(-2M\delta)$. ∎
Now we move on the main Algorithm 8. First, we take some samples to distinguish if $\lambda$ is near 00 or not and if the $r_{i}$’s and the gaps $b_{i}-b_{j}$ are small. If it is near 00 we call $EXP3.P$ otherwise, we try $\tilde{N}(\lambda)$s in the Algorithm 7 until we find the right one.
/* Separate to small and big $\lambda$/
1 for each arm $i\in[K]$ do
       Play arm $i$ for $(T/K)^{2/3}$ rounds. // arm $i$ chosen at random
       /* State now is at: $q_{t}\approx b_{i}$, if $\lambda$ is big enough */
2       for $(T/K)^{2/3}$ rounds do
             Play arm $i$ (same as in Line 1) and update estimator $\hat{\mu}_{i}\leftarrow\hat{r}_{i}+R_{i}/(T/K)^{2/3}$. // $rew_{t}\sim\textsf{Bern}(q_{t}r_{i})$
3
4Choose a random arm $j\in[K]$.
5 for each $i\in[K]\setminus j$ do
6       Play arms $i,j$ alternately for $2(T/K)^{2/3}$ rounds.
7       for $2(T/K)^{2/3}$ rounds do
8             Play arm $i$ and update $\hat{\mu}_{(i,j)}\leftarrow\hat{\mu}_{(i,j)}+R_{i,j}/\sqrt{T}$.
9             Play arm $j$.
10
/* Small $\lambda$. */
11 if $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq 3\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ then
12       Call EXP3.P
13
/* Big $\lambda$. */
14 else
15       $\tilde{N}(\lambda)\leftarrow\log T$
16       Call Algorithm 7 for $\tilde{N}(\lambda)$, arms $i,j$ and get $\hat{\lambda}_{1}$
17       Call Algorithm 7 for $2\tilde{N}(\lambda)$, arms $i,j$ and get $\hat{\lambda}_{2}$
18       $\tilde{N}(\lambda)\leftarrow 4\tilde{N}(\lambda)$
19       while $|\hat{\lambda}_{1}-\hat{\lambda}_{2}|\geq\delta$ do
20             Call Algorithm 7 for $\tilde{N}(\lambda)$, arms $i,j$ and get $\hat{\lambda}_{3}$
21             $\hat{\lambda}_{1}\leftarrow\hat{\lambda}_{2}$
22             $\hat{\lambda}_{2}\leftarrow\hat{\lambda}_{3}$
23             $\tilde{N}(\lambda)\leftarrow 2\tilde{N}(\lambda)$
24      Call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$
ALGORITHM 8 B-DES Unknown $\lambda$
In lemmas E.5 and lemmas E.6 we bound the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$. Then, we bound their difference and prove that if real $\lambda$ is small we call $EXP3.P$ otherwise we are trying to learn real $\lambda$.
Lemma E.5.
For the estimators $\hat{\mu}_{i}$ it holds:
$\Pr\left[|\hat{\mu}_{i}-Y_{i}|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$
@@ 24      Call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$
$Y_{i}=r_{i}\left((1-\lambda)^{(T/K)^{2/3}}\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{T^{2/3}\lambda}(q_{0,i}-b_{i})+b_{i}\right)$
and $q_{0,i}$ the value of state before we start building the estimator.
@@
Let $q_{0,i}$ be the state before we start pulling arm $i$. Then, after $(T/K)^{2/3}$ pulls of this arm the state becomes:
$q_{(T/K)^{2/3}+1,i}$ $=(1-\lambda)^{(T/K)^{2/3}}+\lambda\sum_{s=0}^{T^{2/3}-1}(1-\lambda)^{s}b_{i}$
$=(1-\lambda)^{(T/K)^{2/3}}+\lambda\cdot\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{1-(1-\lambda)}b_{i}$
@@
Then, for another $(T/K)^{2/3}$ rounds the algorithm samples reward from $\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+s,i})$ for $s\in[(T/K)^{2/3}]$, where for each $s\in[(T/K)^{2/3}]$ it applies:
$q_{(T/K)^{2/3}+s,i}$ $=(1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{s,i}-b_{i})+b_{i}$
Thus, the expected reward for each $s\in[(T/K)^{2/3}]$ is $\mathbb{E}[R_{i,s}]=r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$, and
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}q_{t}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
$=r_{i}\left((1-\lambda)^{(T/K)^{2/3}}\frac{1-(1-\lambda)^{(T/K)^{2/3}}}{\lambda}(q_{0,i}-b_{i})+(T/K)^{2/3}b_{i}\right)$
Using Hoeffding’s inequality on $\hat{\mu}_{i}$ we get the result.
@@
For the estimators $\hat{\mu}_{i,j}$ it applies:
$\Pr\left[|\hat{\mu}_{i,j}-Y_{i,j}|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$
@@
$Y_{i,j}=\frac{(1-\lambda)^{2(T/K)^{2/3}}}{(T/K)^{2/3}}\left(r_{i}q_{0,(i,j)}-\frac{r_{i}b_{j}}{2-\lambda}-(1-\lambda)\frac{r_{i}b_{i}}{2-\lambda}\right)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{1-(1-\lambda)^{2}}+\frac{r_{i}b_{j}+(1-\lambda)r_{i}b_{i}}{2-\lambda},$
and $q_{0,(i,j)}$ the value of state before we start building the estimator.
@@
After alternating between a random arm $i$ with arm $j$ for $(T/K)^{2/3}$ the state becomes (lemma E.3):
$q_{2(T/K)^{2/3}+1,(i,j)}=(1-\lambda)^{2(T/K)^{2/3}}q_{0,(i,j)}+\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{2-\lambda}b_{i}.$
And after that for another $(T/K)^{2/3}$ the algorithm samples reward from $\textsf{Bern}(r_{i}\cdot q_{(T/K)^{2/3}+2s+1})$, where
$q_{2(T/K)^{2/3}+2s+1,(i,j)}=(1-\lambda)^{2(T/K)^{2/3}+2s}q_{0,(i,j)}+\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}b_{i}.$
for $s=\{0,...,(T/K)^{2/3}-1\}$. Thus, the expected reward $\mathbb{E}[R_{(i,j),s}]=r_{i}\cdot(1-\lambda)^{2(T/K)^{2/3}+2s}q_{0,(i,j)}$ $+\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{i}$. Thus,
$\mathbb{E}[\sum_{s=0}^{(T/K)^{2/3}-1}R_{(i,j),s}]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{2(T/K)^{2/3}+2s}q_{0,(i,j)}$
$+\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{j}+(1-\lambda)\frac{1-(1-\lambda)^{2(T/K)^{2/3}+2s}}{2-\lambda}r_{i}b_{i}$
$=(1-\lambda)^{2(T/K)^{2/3}}\left(r_{i}q_{0,(i,j)}+\frac{r_{i}b_{j}}{2-\lambda}+(1-\lambda)\frac{r_{i}b_{i}}{2-\lambda}\right)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{1-(1-\lambda)^{2}}$
$+(T/K)^{2/3}\frac{r_{i}b_{j}+(1-\lambda)r_{i}b_{i}}{2-\lambda}.$
Thus, using Hoeffding’s inequality on $\hat{\mu}_{i,j}$ we get the result. ∎
If we take a union bound on all estimators we take that for all estimators $\hat{r}_{i}$ and $\hat{r}_{i,j}$ apply lemma E.5 and lemma E.6 with probability at least $2/T$ as $K<T$. In the continuation of the analysis we assume that applies.
@@
If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$ then the algorithm achieves regret:
@@
If $\lambda$ is in $[0,\tilde{\Theta}(1/T)]$ the states $q_{t}$ during the $K^{1/3}T^{2/3}$ rounds of building the estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$:
$q_{t}\geq(1-\lambda)^{K^{1/3}T^{2/3}}q_{0}\geq(1-1/T)^{K^{1/3}T^{2/3}}>1-\frac{1}{T^{1/3}}$
Thus, $|Y_{i}-Y_{(i,j)}|\leq r_{i}|q_{t}-q_{t}^{\prime}|$, with $t\in[0,K^{1/3}T^{2/3}]$ thus, $|q_{t}-q_{t}^{\prime}|\leq\frac{1}{T^{2/3}}$ and so $|Y_{i}-Y_{(i,j)}|\leq\frac{1}{T^{1/3}}$ for all arms $i,j$ in $[K]$. This means that $|\mu_{i}-\mu_{i,j}|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ and after the $K\cdot(T/K)^{2/3}=K^{1/3}T^{2/3}$ rounds we run $EXP3.P$. From Section 4 we get :
$R_{\textsc{DES}}(T)=2K^{1/3}T^{2/3}+\mathcal{O}(\sqrt{KT\log T})+(1-(1-\lambda)^{T})\textsf{OPT}$
which ends the proof. ∎
@@
If $\max_{i,j\in[K]}\{|b_{i}-b_{j}|\}\leq\sqrt{\log T}/(T/K)^{1/3}$ then $\forall\lambda\in[0,1]$ Algorithm 8 after $K^{1/3}T^{2/3}$ calls $EXP3.P$ and achieves $R_{\textsc{DES}}(T)=\mathcal{O}\left(K^{1/3}T^{2/3}\right)$ regret.